# Large Model Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v1

@@ III-A1 Competition
Competition involves LM agents pursuing their individual objectives, often at the expense of others. This interaction mode is characterized by non-cooperative strategies where agents aim to maximize their own benefits. Non-cooperative game and multi-agent debate have been widely adopted to model the competitive behaviors among LM agents. In non-cooperative games, LM agents engage in strategic decision-making where each LM agent’s goal is to search the Nash equilibrium. LM agents in multi-agent debate [99] involve engaging in structured arguments or debates to defend their positions and challenge the strategies of others.
@@
The cognitive behavior of LLMs, such as self-reflection, has proven effective in solving NLP tasks but can lead to thought degeneration due to biases, rigidity, and limited feedback. Multi-agent debate (MAD) [99] explores interactions among LLMs, where agents engage in a dynamic tit-for-tat, allowing them to correct each other’s biases, overcome resistance to change, and provide mutual feedback. In MAD, diverse role prompts (i.e., personas) are crucial and there are generally three communication strategies: (a) one-by-one debate, (b) simultaneous-talk, and (c) simultaneous-talk-with-summarizer.
@@
III-A2 Partial Cooperation
Partial cooperation occurs when LM agents collaborate to a limited extent, often driven by overlapping but not fully aligned interests [62]. In such scenarios, agents might share certain resources or information while retaining autonomy over other aspects. This interaction mode balances the benefits of cooperation with the need for individual agency and competitive advantage. Partial cooperation can be strategically advantageous in environments where complete cooperation is impractical or undesirable due to conflicting goals or resource constraints. Hierarchical game and coalition formation theory can be employed to model both cooperative and competitive interactions among LM agents.
Hierarchical game [100] structures the interactions of LM agents across different game stages, where they may cooperate at one level and compete at another. Through the coalition formation theory [101], LM agents can form the optimal stable coalition structure under different scenarios and environments, where the optimal stable coalition structure is featured with intra-coalition cooperation and inter-coalition competition. Liu et al. propose DyLAN [102], a multi-layered dynamic LLM agent network for collaborative task-solving, such as reasoning and code generation. DyLAN facilitates multi-round interactions in a dynamic setup with an LLM-empowered ranker to deactivate low-performing agents, an early-stopping mechanism based on Byzantine Consensus to efficiently reach agreement, and an automatic optimizer to select the best agents based on agent importance scores. However, existing cooperative agents rely on learning-based methods, with performance highly dependent on training diversity, limiting their adaptability with unfamiliar teammates, especially in zero-shot coordination. Zhang et al. [85] introduce ProAgent for cooperative agents, featuring four key modules: planner, verificator, controller, and memory. Verificator checks skill feasibility, identifies issues if a skill fails, and triggers re-planning if necessary. If feasible, the controller breaks the skill into executable actions.
III-A3 Cooperation
For connected LM agents, cooperation is a fundamental interaction mode where LM agents work together to achieve common goals, share resources, and optimize collective outcomes. This involves data cooperation, computation cooperation, and information (including knowledge) cooperation.
  • Data Cooperation: LM agents continuously exchange and fuse their individual data (e.g., perceived data) to ensure a comprehensive and up-to-date understanding of their environment, thereby enhancing collective intelligence and enable coordinated actions [81][82].
  • Computation Cooperation: LM agents perform coordinated reasoning, such as service cascades, by distributing computational tasks optimally across agents, leveraging collective processing power to handle complex computations more efficiently [52][85][86][88].
  • Information & Knowledge Cooperation: LM agents share domain-specific knowledge and experiences (e.g., in the format of KGs) to collectively improve their problem-solving capabilities for better-informed actions and decisions, via knowledge bases synchronization and distributed learning algorithms [71][47][83].
The methodologies employed for facilitating cooperation among LM agents include role-playing [103, 90], multi-objective optimization, cooperative game theory [104], Nash bargaining solution, auction mechanism [88], Multi-Agent Reinforcement Learning (MARL) [98], swarm intelligence algorithms, Federated Learning (FL) [86], and the theory of mind [105].
  • Role-playing. Li et al. [30] propose a cooperative agent framework that employs role-playing with inception prompting to guide agents autonomously toward task completion. The system begins with human-provided ideas and role assignments, refined by a task-specifier agent. An AI user and an AI assistant then collaborate through multi-turn dialogues, with the AI user instructing and the assistant responding until the task is completed. ChatDev [90] is a virtual software company operated by “software agents” with diverse roles, e.g., chief officers, programmers, and test engineers. Following the waterfall model, it divides development into design, coding, testing, and documentation, breaking each phase into atomic subtasks managed by a chat chain. MetaGPT [91] integrates human workflows into LLM-based multi-agent collaborations. By encoding Standardized Operating Procedures (SOPs) into prompt sequences, MetaGPT streamlines workflows and reduces errors through agents with human-like expertise. Agents are defined with specific profiles and communicate through structured communication interfaces, such as documents and diagrams, instead of dialogue. Using a publish-subscribe mechanism, agents can freely exchange messages via a shared message pool, publishing their outputs and accessing others’ transparently. Park et al. [103] create a community of 25 generative agents in a sandbox world named Smallville, where agents are represented by sprite avatars. These agents perform daily tasks, form opinions, and interact, mimicking human-like behaviors.
  • Theory of mind. It refers to the ability to understand about others’ hidden mental states, which is essential for social interactions. As LLMs engage more in human interactions, enhancing their social intelligence is vital. Li et al. [105] identify limitations in LLM collaboration and propose a prompt-engineering method to incorporate explicit belief state representations. They also introduce a novel evaluation of LLMs’ high-order theory of mind in teamwork scenarios, emphasizing dynamic belief state evolution and intent communication among agents.
@@ III-A3 Cooperation
In the following, we discuss the detailed cooperation paradigms of LM agents from the perspectives of data cooperation, computation cooperation, and knowledge cooperation.
III-B Data Cooperation for LM Agents
The data cooperation among LM agents involves the modality perspective, the temporal perspective, and the spatial perspective. Effective data cooperation ensures that LM agents can seamlessly integrate and utilize data from diverse sources and modalities, enhancing their capabilities and performance in various applications.
III-B1 Multimodal Data Cooperation
Multimodal data cooperation iemphasizes the fusion of data from various modalities, such as text, images, audio, and video, to provide a comprehensive understanding of the environment. This cooperation allows LM agents to process and interpret information from multiple sources, leading to more accurate and robust decision-making.
  • Data Fusion: By combining data from different modalities, it helps create a unified representation that leverages the strength of each type of data. For example, Wu et al. [106] propose a multi-agent collaborative vehicle detection network named MuDet, which integrates RGB and height map modalities for improved object identification in dense and occluded environments such as post-disaster sites. Gross et al. [81] discuss the use of multimodal data to model communication in artificial social agents, emphasizing the importance of verbal and nonverbal cues for natural human-robot interaction.
  • Cross-Modal Retrieval: By enabling LM agents to retrieve relevant information across different modalities, it enhances their ability to respond to complex queries and scenarios. For example, Gur et al. [107] design an alignment model and retrieval-augmented multi-modal transformers for efficient image-caption retrieval in visual Question-Answering (QA) tasks. By considering the intra-modality similarities in multi-modal video representations, Zolfaghari et al. [72] introduce the contrastive loss in contrastive learning process for enhanced cross-modal embedding. The effectiveness is validated using LSMDC and Youcook2 datasets for video-text retrieval tasks and video captioning tasks. To further enable fine-grained cross-modal retrieval, Chen et al. [108] develop a novel attention mechanism to effectively integrate feature information from different modalities and represent them within a unified space, thereby overcoming semantic gap between multiple data modalities
  • Contextual Understanding: By utilizing multimodal data, it facilitates a richer contextual understanding of the environment with improved accuracy of predictions and actions. For example, Li et al. [78] design a general semantic communication-based multi-agent collaboration framework with enhanced contextual understanding and study a use case in search and rescue tasks.
III-B2 Spatio-temporal Data Cooperation
Spatio-temporal data cooperation for LM agents involves the integration and synchronization of spatial and temporal data across various modalities and sources, enabling LM agents to achieve a comprehensive and dynamic understanding of the environment over time. This cooperation ensures that LM agents can effectively analyze patterns, predict future states, and make informed decisions in real-time, based on both spatial distribution and temporal evolution of data.
  • Spatio-temporal Feature Fusion: Yang et al. [109] introduce SCOPE, a collaborative perception mechanism that enhances spatio-temporal awareness among on-road agents through end-to-end aggregation. SCOPE excels by leveraging temporal semantic cues, integrating spatial information from diverse agents, and adaptively fusing multi-source representations to improve accuracy and robustness. However, [109] mainly works for small-scale scenarios. By capturing both spatial and temporal heterogeneity of citywide traffc, Ji et al. [110] propose a novel spatio-temporal self-supervised learning framework for traffic prediction that improves representation of traffic patterns. This framework uses an integrated module combining temporal and spatial convolutions and employs adaptive augmentation of traffic graph data, supported by two auxiliary self-supervised learning tasks to improve prediction accuracy. To further address data noises, missing information, and distribution heterogeneity in spatio-temporal data, which are overlooked in [109, 110], Zhang et al. [111] devise an automated spatio-temporal graph contrastive learning framework named AutoST. Built on a heterogeneous Graph Neural Network (GNN), AutoST captures multi-view region dependencies and enhances robustness through a spatio-temporal augmentation scheme with a parameterized contrastive view generator.
  • Temporal dynamics with topological reasoning: Chen et al. [82] propose a temporally dynamic multi-agent collaboration framework that organizes agents using directed acyclic graphs to facilitate interactive reasoning, demonstrating superior performance across various network topologies and enabling collaboration among thousands of agents. A key finding in [82] is the discovery of the collaborative scaling law, where solution quality improves in a logistic growth pattern as more agents are added, with collaborative emergence occurring sooner than neural emergence.
III-B3 Space-Air-Ground-Sea Data Cooperation
Space-Air-Ground-Sea (SAGS) data cooperation involves the integration of data from various spatial locations, encompassing satellites, aerial vehicles, ground-based sensors, and maritime sources. This cooperation enables LM agents to have a holistic view of the environment, crucial for applications in areas such as disaster response, environmental monitoring, and logistics. Xu et al. [112] propose an analytical model of coverage probability for ocean Surface Stations (SSs) located far from the coastline. By employing various types of relays including onshore stations, tethered balloons, high-altitude platforms, and satellites, this model establishes communication links between core-connected base stations and SSs via stochastic geometry, adapting relay station selection based on the SS’s distance from the coastline. To further minimize total network latency caused by long-distance transmission and environmental factors, Nguyen et al. [113] broken down the latency optimization problem into three sub-problems: clustering ground users with UAVs, cache placement in UAVs to alleviate backhaul congestion, and power allocation for satellites and UAVs. A distributed optimization approach is devised, which utilizes a non-cooperative game for clustering, a genetic algorithm for cache placement, and a quick estimation technique for power allocation.
Figure 11: The taxonomy of computation cooperation of LM agents.
III-C Computation Cooperation for LM Agents
The computation cooperation paradigm of LM agents includes horizontal/vertical cooperation and cross-layer cooperation.
① Horizontal/vertical LM collaboration. It can be classified into three modes: horizontal cooperation, vertical cooperation, and hybrid cooperation.
@@ III-C Computation Cooperation for LM Agents
Figure 12: Illustration of horizontal collaboration.
1) Horizontal collaboration refers to multiple LM agents completing the same task independently at the same time, and then summarizing and integrating their respective outputs to generate the final result. It allows the collaborative system to scale horizontally by adding more LM agents to handle complex and dynamic tasks. The parallel processing and multi-angle perspectives contribute to increased robustness and diversity, minimizing errors and biases that could arise from a single model’s limitations. As depicted in Fig. 12, each LM agent independently analyzes different aspects of stock including trend, news, and industry to determine whether it is a good investment. Horizontal collaboration has been applied in various LM agent systems. For example, ProAgent [85] designs a framework where LM agents work as teammates, analyzing each other’s intentions and updating their own beliefs based on the observed behaviors of their peers. This enhances collective decision-making by allowing agents to adapt dynamically in real time. Similarly, DyLAN [114] assembles a team of strategic agents that communicate through task-specific queries in a dynamic interaction architecture, enabling multiple rounds of interaction to improve both efficiency and overall performance. However, the independent analysis of each LM agent may lead to inconsistent outputs, making it difficult to aggregate the final conclusions. Therefore, horizontal collaboration requires effective mechanisms to handle disagreements between LM agents and ensure that their contributions complement each other, which may increase the complexity of the coordination process.
Figure 13: Illustration of vertical collaboration.
2) Vertical collaboration is to decompose complex tasks into multiple stages, and different LM agents complete the tasks of each stage in turn. After each agent completes the task of its own stage, it passes the result to the next agent until the entire task is successfully solved. As depicted in Fig. 13, for medical image analysis, the first agent may extract basic visual features, the second agent analyzes these features in more depth, and the final agent produces a diagnosis. This sequential process enables a step-by-step refinement, allowing complex problems to be broken down and tackled more efficiently by leveraging specialized agents at each stage. For example, Jiang et al. [28] present CommLLM, a multi-agent system for natural language-based communication tasks. It features three components: 1) multi-agent data retrieval, which uses condensate and inference agents to refine 6G communication knowledge; 2) multi-agent collaborative planning, employing various planning agents to develop solutions from multiple viewpoints; and 3) multi-agent reflection, which evaluates solutions and provides improvement recommendations through reflection and refinement agents. However, a limitation of vertical collaboration is that higher-level LM agents rely on the accuracy of lower-level outputs, making the process vulnerable to error propagation. Mistakes made early can compromise the final outcome, therefore this approach demands high precision from each agent to ensure overall system reliability.
3) Hybrid collaboration. In practical LM agent environments, real-world applications often involve both horizontal and vertical collaboration paradigms, resulting in the hybrid collaboration. For example, in addressing highly complex tasks, the problem is first broken down into manageable sub-tasks, each assigned to specialized LM agents. Here, horizontal collaboration enables agents to perform parallel evaluations or validations, while vertical collaboration ensures that sequential processing stages refine the task further. The computational collaboration among LM agents involves a mix of horizontal and vertical approaches, such as coordinating parallel assessments across agents while sequentially integrating outputs through defined stages, thus optimizing task execution and enhancing overall system performance.
@@
② Cross-layer computation cooperation. For cross-layer computation cooperation among LM agents, it can be classified into two modes: cloud-edge cooperation and edge-edge cooperation.
TABLE VI: A Summary of Cloud-edge Computation Cooperation Approaches for LM Agents
Strategy Feature Technology Ref.
Transfer learning or compression Reduction of training costs for multimodal LLMs Transfer of VPG across LLMs [115]
Task-agnostic structural pruning for LLMs LoRA for recovery after pruning with minimal data [116]
Caching KV cache compression in LLM inference Selective storage of pivotal tokens [117]
Joint caching & inference for LLM agents Cached model-as-a-resource, age-of-thought [88]
Hardware Acceleration Efficient LLM inference on edge devices SIMD-based 4-bit multiplier, TRIP matrix multiplication [118]
Sharding Collaborative inference of LLMs on edge and cloud Model partitioning and adaptive algorithm for optimization [119]
1) Cloud-Edge Cooperation. In real-world scenarios, running a complete LM agent requires substantial computational resources. Edge servers, often equipped with fewer devices and resources than cloud servers, typically lack the capacity to support the operation of a complete LM agent. The cloud-edge collaboration approach enables edge servers to support LM agents effectively. A range of strategies have been developed to optimize resource utilization, enhance performance, and reduce operational costs, encompassing transfer learning, model compression, caching, hardware acceleration, and model sharding. In the literature, there exist the following types of cloud-edge cooperation paradigms: orchestration between LM and Smaller Models (SMs), lightweight edge LM deployment, sharding-based edge LM inference, and cloud-edge LM service optimizations.
  • LM-SM orchestration via transfer learning or compression: It deploys a complete LM agent in the cloud and create smaller, specialized LM agents through transfer learning or model compression. The agents run on smaller models with reduced parameter scales, are then deployed on edge servers, requiring fewer resources while maintaining high performance. VPGTrans [115] addresses the high computational costs of training visual prompt generators for multimodal LLMs by significantly reducing GPU hours and training data needed via transferring an existing visual prompt generator from one multimodal LLM to another. LLM-Pruner [116] compresses LLMs without compromising their multi-task solving and language generation abilities, using a task-agnostic structural pruning method that significantly reduces the need for extensive retraining data and computational resources.
  • Cloud-Edge LM deployment via quantization and hardware acceleration: It employs specialized hardware to accelerate the deployment and operation of LMs at the edge. Hardware accelerators, such as GPUs and TPUs, can enhance computational efficiency and provide robust support for cloud training and edge deployment. Agile-Quant [118] enhances the efficiency of LLMs on edge devices through activation-guided quantization and hardware acceleration by utilizing a SIMD-based 4-bit multiplier and efficient TRIP matrix multiplication, achieving up to 2.55x speedup while maintaining high task performance.
  • Cloud-Edge LM deployment via Split Learning (SL): SL is an emerging distributed ML paradigm in which the model is split into several parts [120]. SL is typically implemented in three settings: two-part single-client, two-part multi-client, and U-Shape configurations [121]. The goal of SL is to offload complex computations and enhance data privacy by keeping the preceding model segments local. This approach addresses the substantial computational demands and privacy concerns associated with training and inference in LM agents. For example, Xu et al. [120] propose a cloud-edge-end computing system for LLM agents using SL, where mobile agents with local models (0-10B parameters) handle real-time tasks, and edge agents with larger models (over 10B parameters) provide complex support by incorporating broader contextual data. They also study a real case, where mobile agents create localized accident scene descriptions, which are then enhanced by edge agents to generate comprehensive accident reports and actionable plans.
  • Cloud-Edge LM inference via model sharding: It distributes LM shards across heterogeneous edge servers to accommodate varying device capabilities and resource conditions. By partitioning the LM into smaller shards, it allows that each edge server only handles a manageable portion of the LM, optimizing resource utilization and performance. For example, EdgeShard [119] improves LLM inference by partitioning LMs into shards for deployment on distributed edge devices and cloud servers, optimizing latency and throughput with an adaptive algorithm, achieving up to 50% latency reduction and 2x throughput improvement.
  • Cloud-Edge LM services via caching and resource optimization: It utilizes caching mechanisms and other resource optimization techniques to configure and support LM agents on edge servers. By caching frequently used data and intermediate results, the system reduces the need for continuous high-volume data transfer between the cloud and the edge node. Scissorhands [117] reduces the memory usage of the KV cache in LLMs during inference by selectively storing pivotal tokens, thus maintaining high throughput and model quality without requiring model finetuning. Xu et al. [88] propose a joint caching and inference framework for edge intelligence in space-air-ground integrated networks to efficiently deploy LLM agents. They design a new cached model-as-a-resource paradigm and introduce the concept of age-of-thought for optimization, along with a deep Q-network-based auction mechanism to incentivize network operators.
3) Edge-Edge Cooperation. Edge-edge cooperation involves the collaboration between multiple edge devices to enhance computational efficiency and resource utilization. In practical scenarios, edge servers need to protect local data privacy while lacking the resources to train LM agents independently. By leveraging FL, edge-edge cooperation enables decentralized training of LM agents across edge devices without requiring data to be centralized, thus preserving data privacy and reducing latency. For example, to tackle issues of data scarcity and privacy in LLM development, Chen et al. [86] propose federated LLM, combining federated pre-training, fine-tuning, and prompt engineering to enable effective collaborative training without compromising data privacy. By utilizing FL, edge-edge cooperation also ensures robust and scalable deployment of LM agents in resource-constrained environments, allowing edge devices to process and analyze data locally while maintaining privacy and efficiency.
III-D Knowledge Cooperation for LM Agents
@@
Knowledge can be divided into explicit knowledge and implicit knowledge. Explicit knowledge is the information directly utilized by the AI model, such as external databases and KGs. Implicit knowledge is obtained through the optimization of AI model’s internal parameters (i.e., weights and biases) [122]. As depicted in Fig. 14, the knowledge cooperation of LM agents generally involves three aspects: knowledge synchronization, knowledge searching, as well as the cooperation between knowledge and LMs.
Figure 14: Illustration of knowledge synchronization in LM agents. (a) Parameterized knowledge transfer in knowledge synchronization, where parameter knowledge extracted from the teacher model is injected into the student model and further trained with high-quality small sample data to obtain the target network. (b) Retrieval-enhanced knowledge search using KGs, where entities detected in the query are used to retrieve corresponding triples from KGs to assist LMs in reasoning. (c) KG completion by orchestrating knowledge and LMs, where five example triples and the missing triple are used as queries for LMs to complete entities, which are then fed back into both KGs and LMs.
TABLE VII: A Summary of Knowledge Cooperation Approaches for LM Agents
Mode Type Technology Ref.
@@
Online Distillation LMs generate high-quality reasoning processes for models to learn from [123]
Knowledge Transfer Offline Distillation LMs generate distilled data or fine-tuning instructions to train models [124]
@@
Parametric Knowledge Transfer Extract and align the knowledge parameters of LMs to models [125]
@@
Knowledge fusion Train target model via predictive probability distributions of LMs [74]

### v2

@@ III-E2 Distributed Competition
Competition involves LM agents pursuing their individual objectives, often at the expense of others. This interaction mode is characterized by non-cooperative strategies where agents aim to maximize their own benefits. Multi-agent debate (MAD) and non-cooperative game have been widely adopted to model the competitive behaviors between autonomous agents.
@@
  • MAD. The cognitive behavior of LLMs, such as self-reflection, has proven effective in solving NLP tasks but can lead to thought degeneration due to biases, rigidity, and limited feedback. LM agents in MAD [104] involve engaging in structured arguments or debates to defend their positions and challenge the strategies of others. MAD explores interactions among LLMs, where agents engage in a dynamic tit-for-tat, allowing them to correct each other’s biases, overcome resistance to change, and provide mutual feedback. In MAD, diverse role prompts (i.e., personas) are crucial and there are generally three communication strategies: (a) one-by-one debate, (b) simultaneous-talk, and (c) simultaneous-talk-with-summarizer, as illustrated in Fig. 17. In one-by-one debate strategy, debater agents respond sequentially in a fixed order, generating replies based on the ongoing discussion. Each agent’s chat history is updated with previous responses to maintain context. In simultaneous-talk debates, debater agents respond asynchronously within each round, mitigating the influence of speaking order on the debate’s flow. The simultaneous-talk-with-summarizer debate strategy enhances simultaneous-talk by incorporating a summarizer LLM. After each round, the summarizer condenses the discussion into a cohesive summary, which replaces the chat histories of all agents, ensuring shared context.
  • Non-cooperative Game. In non-cooperative games, LM agents engage in strategic decision-making where each LM agent’s goal is to seek the Nash equilibrium. Various non-cooperative games can be employed such as Nash bargaining, auctions, matching games. For instance, Xu et al. [105] design a new cached model-as-a-resource paradigm and introduce the concept of age-of-thought to efficiently deploy LLM agents. They also devise a deep Q-network-based auction mechanism to incentivize network operators.
III-E3 Partial Cooperation
Partial cooperation occurs when LM agents collaborate to a limited extent, often driven by overlapping but not fully aligned interests [63]. In such scenarios, agents might share certain resources or information while retaining autonomy over other aspects. This interaction mode balances the benefits of cooperation with the need for individual agency and competitive advantage. Partial cooperation can be strategically advantageous in environments where complete cooperation is impractical or undesirable due to conflicting goals or resource constraints. For instance, DyLAN [106] is a multi-layered LLM agent network for collaborative task-solving with teammates, such as reasoning and code generation. DyLAN facilitates multi-round interactions in a dynamic setup with an LLM-empowered ranker to deactivate low-performing agents, an early-stopping mechanism based on Byzantine consensus to efficiently reach agreement, and an automatic optimizer to select the best agents based on agent importance scores.
Hierarchical game and coalition formation theory can be employed to model partially cooperative interactions among LM agents. Hierarchical game [107] structures the interactions of LM agents across different game stages, where they may cooperate at one level and compete at another. Through the coalition formation theory [83], LM agents can form the optimal stable coalition structure under different scenarios and environments, where the optimal stable coalition structure is featured with intra-coalition cooperation and inter-coalition competition.
III-E4 Lessons Learned
In distributed collaboration, SL and MoE focus on model segmentation, while FL and role-playing emphasize model integration. Within the overall cloud-edge collaboration for LM agent systems, key challenges emerge in the joint consideration of LM segmentation (e.g., SL and MoE), LM placement, LM replacement, LM routing, and knowledge caching, alongside the inherent heterogeneity and dynamics of edge environments. Effective deployment and operation of LM agents across cloud and edge infrastructures require a delicate balance between LM segmentation and LM integration strategies, with an emphasis on optimizing resource allocation and minimizing service latency. Additionally, solutions should be adaptable to the constantly changing network conditions and hardware limitations inherent to edge devices. Collaborative models such as SL and MoE can significantly reduce computational burdens on edge devices, while federated strategies and role-playing can enhance cooperation among distributed agents. Future research should explore the seamless integration of these techniques, taking into account both static and dynamic aspects of edge networks, to provide more efficient and scalable solutions for large-scale applications.
@@ III-E4 Lessons Learned
III-F Cooperation Paradigms of Connected LM Agents
@@
As illustrated in Fig. 18, the detailed cooperation paradigms of LM agents under cloud-edge-end architecture involves three perspectives: data cooperation, computation cooperation, and knowledge cooperation.
  • Data Cooperation: Within a common task, LM agents continuously exchange and fuse their individual data (e.g., task-oriented sensory data) to ensure a comprehensive and up-to-date understanding of their task environment, thereby enhancing collective intelligence and enable coordinated actions.
  • Computation Cooperation: LM agents perform coordinated reasoning, such as service cascades, by distributing computational tasks optimally across agents, leveraging collective processing power to handle complex computations more efficiently.
  • Knowledge Cooperation: LM agents share domain-specific knowledge and experiences (e.g., in the format of KGs) to collectively improve problem-solving capabilities for better-informed actions and decisions, via knowledge synchronization, fusion, and retrieval and distributed learning algorithms.
@@
Figure 19: Illustration of (a) horizontal collaboration and (b) vertical collaboration paradigms for multiple LM agents.
1) Data Cooperation for LM Agents: The data cooperation among LM agents involves the modality perspective and the spatio-temporal perspective.
a) Multimodal Data Cooperation emphasizes the fusion of data from various modalities, such as text, images, audio, and video, to provide a comprehensive understanding of the environment. This cooperation allows LM agents to process and interpret information from multiple sources, leading to more accurate and robust decision-making. (i) By combining data from different modalities, it helps to create a unified representation that takes advantage of the strengths of each type of data. For instance, Gross et al. [108] discuss the use of multimodal data to model communication in artificial social agents, emphasizing the importance of verbal and nonverbal cues for natural human-robot interaction. (ii) By enabling LM agents to retrieve relevant information across different modalities, it enhances their ability to respond to complex queries and scenarios. For instance, by considering the intra-modality similarities in multi-modal video representations, Zolfaghari et al. [109] introduce the contrastive loss in contrastive learning process for enhanced cross-modal embedding. The effectiveness is validated using LSMDC and Youcook2 datasets for video-text retrieval task and video captioning task.
b) Spatio-temporal Data Cooperation involves the integration and synchronization of spatial and temporal data across various modalities and sources, enabling LM agents to achieve a comprehensive and dynamic understanding of the environment over time. This cooperation ensures that LM agents can effectively analyze patterns, predict future states, and make informed decisions in real-time, based on both spatial distribution and temporal evolution of data. Yang et al. [86] introduce SCOPE, a collaborative perception mechanism that enhances spatio-temporal awareness among on-road agents through end-to-end aggregation. SCOPE excels by leveraging temporal semantic cues, integrating spatial information from diverse agents, and adaptively fusing multi-source representations to improve accuracy and robustness. However, [86] mainly works for small-scale scenarios. By capturing both spatial and temporal heterogeneity of citywide traffc, Ji et al. [110] propose a novel spatio-temporal self-supervised learning framework for traffic prediction that improves representation of traffic patterns. This framework uses an integrated module combining temporal and spatial convolutions and employs adaptive augmentation of traffic graph data, supported by two auxiliary self-supervised learning tasks to improve prediction accuracy.
@@
2) Computation Cooperation for LM Agents: It can be classified into three modes: horizontal cooperation, vertical cooperation, and hybrid cooperation.
a) Horizontal collaboration refers to decompose complex tasks into manageable sub-tasks, with multiple LM agents independently completing their assigned tasks in parallel. The results are then summarized and integrated to generate the final outcome. This approach enables the collaborative system to scale horizontally by adding more LM agents, allowing for more complex and dynamic tasks to be handled. The parallel processing and multi-angle perspectives contribute to increased robustness and reduced errors and biases that may arise from relying on a single agent. As shown in Fig. 19(a), each LM agent independently analyzes different aspects of stock performance such as trend, news, and industry to assess its investment potential. Horizontal collaboration has been successfully implemented in various LM agent systems. For instance, ProAgent [84] employs a framework where LM agents collaborate as teammates, analyzing each other’s intentions and updating their beliefs based on observed behaviors of their peers. This enhances collective decision-making by allowing agents to adapt dynamically in real time. Similarly, DyLAN [106] assembles a team of strategic agents that communicate through task-specific queries, allowing multiple rounds of interaction to enhance both efficiency and overall performance. However, the independent analysis of each LM agent may lead to inconsistent outputs, complicating the aggregation of the final conclusions. Therefore, effective mechanisms are needed to address disagreements and ensure that the agents’ contributions are complementary, which may increase the complexity of the coordination process.
b) Vertical collaboration involves decomposing complex tasks into multiple stages, with different LM agents handling each stage sequentially. After completing their respective tasks, each agent passes the result to the next agent, until the entire task is successfully completed. As illustrated in Fig. 19(b), in medical image analysis, the first agent might extract basic visual features, the second agent could analyze these features in greater depth, and the final agent produces a diagnosis. This sequential approach enables a step-by-step refinement, allowing complex problems to be broken down and tackled more efficiently by leveraging specialized agents at each stage. For instance, Jiang et al. [24] present CommLLM, a multi-agent system for natural language-based communication tasks. It comprises three components: 1) multi-agent data retrieval, which uses condensate and inference agents to refine 6G communication knowledge; 2) multi-agent collaborative planning, employing various planning agents to generate solutions from multiple viewpoints; and 3) multi-agent reflection, which evaluates solutions and suggests improvements through reflexion and refinement agents. However, a limitation of vertical collaboration is that higher-level LM agents rely on the accuracy of lower-level outputs, making the process vulnerable to error propagation. Mistakes made early in the sequence can compromise the final outcome, highlighting the need for high precision at each stage to maintain the overall system’s reliability.
c) Hybrid collaboration. In practical LM agent environments, real-world applications often require a combination of horizontal and vertical collaboration, resulting in hybrid collaboration, as illustrated in Fig. 16. For instance, when addressing highly complex tasks, the problem is first broken down into manageable sub-tasks, each assigned to specialized LM agents. Horizontal collaboration allows agents to perform parallel evaluations or validations, while vertical collaboration ensures that the task is refined through sequential processing stages. This computational collaboration blends both paradigms, coordinating parallel assessments across agents while sequentially integrating their outputs through defined stages, thus optimizing task execution and enhancing overall system performance. For instance, Chen et al. [111] propose a dynamic multi-agent collaboration framework that organizes agents using directed acyclic graphs to facilitate interactive reasoning. Their framework demonstrates superior performance across various network topologies and enables collaboration among thousands of agents. A key finding in [111] is the discovery of the collaborative scaling law, where solution quality improves in a logistic growth pattern as more agents are added, with collaborative emergence occurring faster than neural emergence.
@@
3) Knowledge Cooperation for LM Agents: Knowledge can be broadly categorized into explicit knowledge and implicit knowledge. Explicit knowledge refers to structured, codified, and easily accessible information that can be articulated, documented, and shared, such as external databases and KGs. Implicit knowledge is embedded within the model and arises through the optimization of its internal parameters, such as weights and biases, during training or fine-tuning [112]. The knowledge cooperation generally involves three consecutive aspects: knowledge synchronization, knowledge fusion, and knowledge retrieval.
a) Knowledge sharing. It includes the sharing and updating of knowledge between multiple LM agents to ensure consistency in decision making.
① Knowledge transfer is a common method of knowledge sharing that transfers the implicit knowledge, in the form of learned parameters, from one LM agent to another. Kang et al. [113] propose an online distillation method that enhances LLMs by retrieving relevant knowledge from external knowledge bases, to generate high-quality reasoning processes. The knowledge (e.g., reasoning results) of small language models can be leveraged to improve the performance of LLMs in knowledge-intensive tasks. However, as the scale of LLMs grows, LLM training via such knowledge transfer becomes complex and computationally intensive. To address this issue, Zhong et al. [114] propose a parametric knowledge transfer method that extracts and aligns knowledge parameters using sensitivity techniques and uses the LoRA module as an intermediary mechanism to inject this knowledge into smaller models, transferring the implicit knowledge of smaller models.
② Knowledge update. Knowledge alignment is a prerequisite to update knowledge among LM agents. Zhang et al. [115] propose a fully automatic KG alignment method, using LLMs to identify and align entities in different KGs, aiming to solve the heterogeneity problem between different KGs and integrate multi-source data. (i) For explicit knowledge updates, Tandon et al. [116] pair LMs with a growing memory to train a correction model, where users identify output errors and provide general feedbacks on how to correct them. (ii) For implicit knowledge updates, continual learning ensures that AI model can continuously learn while receiving new tasks and new data without forgetting previously learned knowledge. Qin et al. [117] propose ELLE, which flexibly extends the breadth and depth of existing PLMs, allowing the model to continuously grow as new data flows in.
b) Knowledge fusion. It contains the knowledge fusion and completion phases between multiple LM agents.
① Knowledge integration, which integrates and optimizes knowledge from different LM agents to form a robust and comprehensive public knowledge base for a common task. Jiang et al. [118] propose an ensemble framework named LLM-blender that directly aggregates the outputs of different models to improve prediction performance and robustness. However, it requires maintaining multiple trained LLMs and executing each LLM during inference, which can be impractical for LM agents. To address this issue, Wan et al. [79] propose an implicit knowledge fusion framework that evaluates the predictive probability distributions of multiple LLMs and uses the distribution for continuous training the target model.
② Knowledge completion, which infers missing facts in a given KG. Compared with traditional KG completion methods, LLMs hold the potential to enhance KG completion performance via encoding text or generating facts. Shen et al. [119] use LLMs as encoders, primarily capturing the semantic information of KG triples through the model’s forward pass, and then reconstructing the KG’s structure by calculating a loss function, thereby better integrating semantic and structural information. Apart from KG completion, LLMs can be employed to verify KGs. Han et al. [120] propose a prompt framework for iterative KG verification, using small LLMs to correct errors in KG generated by LLMs such as ChatGPT.
TABLE VI: A Summary of Knowledge Cooperation Approaches for LM Agents
Stage Type Technology Ref.
@@
Online knowledge distillation LMs generate high-quality reasoning processes for models to learn. [113]
@@
Offline knowledge distillation Extract and align parameters of LLMs to smaller models. [114]
Knowledge Sharing Explicit knowledge update Adjust LM via knowledge bases and feedbacks. [116]
Implicit knowledge update Selective fine-tuning of model’s internal parameter knowledge. [117]
@@
Implicit knowledge fusion Train target model via predictive probability distributions of LMs. [79]